OpenStudy (anonymous):

Selecting a new health club. The one he likes has monthly dues of $25 and the start-up fee of $300. How many months will he need to belong to have spent $1,000

OpenStudy (anonymous):

700/25 = 28 months

hero (hero):

1000 = 300 + 25x 700 = 25x 700/25 = x 28 = x

hero (hero):

Prometheus 69 took a short cut